# Optimal Liouville theorems for the Lane-Emden equation on Riemannian manifolds

### v1

@@ 4 Gradient estimates
Let $M^{n}$ be a complete and $n$-dimensional Riemannian manifold with the Ricci curvature bounded from below by a nonpositive constant $-(n-1)\kappa$. Assume for some constant $\delta_{0}<1$,
@@
Let $M^{n}$ be a complete and $n$-dimensional Riemannian manifold with the Ricci curvature bounded from below by a nonpositive constant $-(n-1)\kappa$. Assume $p>1$ and the condition (2.3) holds. If $u$ is a positive and weak solution to (2.1), then the following pointwise inequality holds for every $\lambda\geq\lambda_{0}\coloneqq\frac{p}{1-\delta_{0}^{+}}$
@@
It is well known that $u$ is $C^{2}$ in $\left\{\nabla u\neq 0\right\}$ (see [14, 45]). We consider new functions
@@
$\left\langle A_{w}\left(\nabla F\right),\nabla F\right\rangle=p^{2}(p-1)w_{1}^{2p-2}\left((p-1)w_{11}^{2}+\sum_{j=2}^{n}w_{1j}^{2}\right),$
@@
$\left\langle A_{w}\left(\nabla F\right),\mathbf{W}\right\rangle=p(p-1)w_{1}^{2p-2}w_{11},$
we get
@@
$-\dfrac{(n-3)p+2}{(n-1)p}\left\langle A_{w}\left(\nabla F\right),\mathbf{W}\right\rangle+\left(\dfrac{(n+1)(p-1)}{n-1}u^{1-p}f(u)-u^{2-p}f^{\prime}(u)\right)F-(n-1)\kappa F^{2-2/p}$
$\geq$ $\left(1-\delta\right)\dfrac{(p-1)^{2}F^{2}}{n-1}+\left(1-\delta\right)\dfrac{u^{2-2p}f(u)^{2}}{n-1}+\dfrac{2\delta(p-1)}{n-1}u^{1-p}\left\lvert f(u)\right\rvert F$
@@
for every $\delta\in[0,1)$. Thus, for $\delta=\delta_{0}^{+}=\max\left\{0,\delta_{0}\right\}\in[0,1)$, we have
$\dfrac{(n(1-\delta_{0}^{+})-1)(p-1)}{(1-\delta_{0}^{+})(n-1)p^{2}}=\dfrac{((n-1)(1-\delta_{0}^{+})-\delta_{0}^{+})(p-1)}{(1-\delta_{0}^{+})(n-1)p^{2}}\geq-\dfrac{\delta_{0}^{+}}{(1-\delta_{0}^{+})p},\quad\quad\dfrac{\left\lvert(n-3)p+2\right\rvert}{(n-1)p}\leq 1,$
and conclude
@@
For every real number $\lambda>0$, we have from (4.4)
@@
Let $M^{n}$ be a complete and $n$-dimensional Riemannian manifold with the Ricci curvature bounded from below by a nonpositive constant $-(n-1)\kappa$. Assume for some constant $\delta_{0}<1$,
@@
Here we used the following well known fact about the regularity of $u$ (see [45]):
$u\in C_{loc}^{2,\beta}\left(M\setminus Z\right)\cap W^{2,2}_{loc}\left(M\right)\cap C^{1,\beta}_{loc}\left(M\right)$
and for each compact domain $\Omega\Subset M$
$\left\lvert\nabla u\right\rvert^{p-2}\nabla^{2}u\in L^{2}\left(\Omega\setminus Z\right)$
where $Z=\left\{\nabla u\neq 0\right\}$ is the critical set of $u$.
@@
On the other hand, integration by parts, we get
@@
$-\dfrac{4}{p+\lambda_{0}+t-2}\int_{M}\left\langle A_{w}\left(H^{\frac{p+\lambda_{0}+t-2}{2}}\right),\nabla\eta\right\rangle H^{\frac{p+\lambda_{0}+t-2}{2}}\eta$
@@
$\left(\dfrac{t+1}{t},\,t\right),$
@@
$\leq$ $\dfrac{2p(t+1)^{2}}{t}\left(\int_{B_{R}}H^{2t+2}\eta^{2t+2}\right)^{\frac{t}{t+1}}\left(\int_{B_{R}}\left\lvert\nabla\eta\right\rvert^{2t+2}\right)^{\frac{1}{t+1}}+2(n-1)\kappa\left(\int_{B_{R}}H^{2t+2}\eta^{2t+2}\right)^{\frac{t}{t+1}}\left(\int_{B_{R}}\eta^{2t+2}\right)^{\frac{t}{t+1}},$
@@
holds for all cutoff function $\eta$. Fixed $q_{0}$ as in 4.2 and set
@@
$\geq$ $\tau H^{(\tau-1)\lambda_{0}}\lambda_{0}\left(\dfrac{(p-1)^{2}\left(1-\delta_{0}^{+}\right)}{n-1}H^{\lambda_{0}+p}-(n-1)\kappa H^{\lambda_{0}+p-2}-p(p-1)H^{\lambda_{0}+p-2}\left\lvert\nabla H\right\rvert\right)$
@@
Since the Ricci curvature is bounded from below by $-(n-1)\kappa$, applying the Bishop-Gromov volume comparison theorem, we get
@@ 5 Liouville theorems
We begin with the following simple case.
@@
Let $u$ be a positive and weak solution to (2.1). According to (3.10) in 3.5, under the assumption, the following inequality holds in $\left\{\nabla u\neq 0\right\}$
$\operatorname{div}\left(u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\nabla_{\mathbf{X}}\mathbf{X}-\dfrac{\operatorname{div}\mathbf{X}}{n}\mathbf{X}\right)\right)\geq 0.$
The equality holds only if
$\alpha\left(1-\dfrac{(n-p)\alpha}{(n+1)p-n}\right)=0,\quad f(u)=Au^{\alpha},\quad Ric\left(\nabla u,\nabla u\right)=0,$
whenever $\nabla u\neq 0$ and $A$ is a constant. Thus, $u$ must be a constant function. ∎
Another application of 3.5 on compact manifold is that we provide a new and simple proof of Bidaut-Veron and Veron’s Liouville result [5, Theorem 6.1] on compact manifolds.
Theorem 5.2 ([5, Theorem 6.1]).
Let $(M,g)$ be a closed manifold with dimension $n\geq 2$ and $u$ be a positive solution of
@@
$0=$ $\int_{M}u^{a}\mathrm{tr}(\mathscr{X}^{2})+u^{a+2b}\mathrm{Ric}\left(\nabla u,\nabla u\right)$
@@
$0=$ $\int_{M}u^{\frac{2(n-1)}{n+2}q}\mathrm{tr}(\mathscr{X}^{2})+\int_{M}u^{\frac{2(n-3)}{n+2}q}\left(\mathrm{Ric}\left(\nabla u,\nabla u\right)-\dfrac{n-1}{n}\left(q-1\right)\lambda\left\lvert\nabla u\right\rvert^{2}\right)$
$+\int_{M}\frac{(n-1)}{n+2}q\left(1-\frac{n-2}{n+2}q\right)u^{\frac{2(n-3)}{n+2}q-2}\left\lvert\nabla u\right\rvert^{4}.$
@@ 1. Case 1)
    If $q<\frac{n+2}{n-2}$, then it is easy to see that
    $0\geq\int_{M}\frac{(n-1)}{n+2}q\left(1-\frac{n-2}{n+2}q\right)u^{\frac{2(n-3)}{n+2}q-2}\left\lvert\nabla u\right\rvert^{4}$
    implies $\nabla u=0$ and $u$ is a constant.
@@ 2. Case 2)
    If $q=\frac{n+2}{n-2},(M,g)=(\mathbb{S}^{n},g_{0})$ and (5.1) is strict, then
    $0\geq\int_{M}u^{\frac{2(n-3)}{n-2}}\left(\mathrm{Ric}\left(\nabla u,\nabla u\right)-\dfrac{n-1}{n}\left(q-1\right)\lambda\left\lvert\nabla u\right\rvert^{2}\right)$
    implies $\nabla u=0$ and $u$ is a constant.
@@ 3. Case 3)
    If $q=\frac{n+2}{n-2}$, $(M,g)\neq(\mathbb{S}^{n},g_{0})$ and (5.1) is not strict, then we have
    $\int_{M}u^{\frac{2(n-1)}{n-2}}\mathrm{tr}(\mathscr{X}^{2})=0.$
    Hence, the fact $u>0$ forces $\mathrm{tr}(\mathscr{X}^{2})=0$, i.e.,
    $\mathrm{tr}(\mathcal{X}-\frac{\mathrm{tr}\mathcal{X}}{n})^{2}=0.$
    Noting
    $\mathcal{X}=\nabla(u^{b}\nabla u)=\nabla(u^{-\frac{2}{n-2}}\nabla u)=\begin{cases}\frac{n-2}{n-4}\nabla\nabla(u^{\frac{n-4}{n-2}}),&\text{ if }n\neq 4;\\ \nabla\nabla\ln u,&\text{ if }n=4,\end{cases}$
    hence, we have
    $\nabla\nabla(u^{\frac{n-4}{n-2}})=\frac{1}{n}\Delta(u^{\frac{n-4}{n-2}})g,\quad\text{ if }n\neq 4;$
@@
    According to the rigidity theorem [56, Theorem E] (see [27] for more detailed discussions), the equation
    $\nabla\nabla f=\frac{1}{n}\Delta fg$
    admits a non-trivial solution if and only if $(M,g)=(\mathbb{S}^{n},g_{0})$. Thus $u$ must be a constant.
@@
It follows from (3.10) in 3.5, the following inequality holds
$\begin{split}\operatorname{div}\left(u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\mathscr{X}(\mathbf{X})\right)\geq\,&u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\mathrm{tr}(\mathscr{X}^{2})\right)+A_{n,p,\alpha}u^{-\frac{((n+1)p-2n)\alpha}{(n+1)p-n}-2}\left\lvert\nabla u\right\rvert^{2p}\\ &+\dfrac{n-1}{n}\left(\alpha f(u)-uf^{\prime}(u)\right)u^{-\frac{((n+1)p-2n)\alpha}{(n+1)p-n}-1}\left\lvert\nabla u\right\rvert^{p}\end{split}$ (5.4)
@@
$\int_{M}\operatorname{div}\left(u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\mathscr{X}(\mathbf{X})\right)\eta^{\gamma}$
$=$ $-\gamma\int_{M}u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left\langle\mathscr{X}(\mathbf{X}),\nabla\eta\right\rangle\eta^{\gamma-1}$
$\leq$ $\gamma\int_{M}u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left\lvert\mathscr{X}(\mathbf{X})\right\rvert\left\lvert\nabla\eta\right\rvert\eta^{\gamma-1}$
$\leq$ $\sqrt{\frac{1+(p-1)^{2}}{2(p-1)}}\gamma\int_{M}u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\mathrm{tr}\left(\mathscr{X}^{2}\right)\right)^{1/2}u^{-\frac{n(p-1)\alpha}{(n+1)p-n}}\left\lvert\nabla u\right\rvert^{p-1}\left\lvert\nabla\eta\right\rvert\eta^{\gamma-1}$
$\leq$ $\sqrt{\frac{1+(p-1)^{2}}{2(p-1)}}\gamma\left(\int_{M}u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\mathrm{tr}\left(\mathscr{X}^{2}\right)\right)\eta^{\gamma}\right)^{1/2}\left(\int_{M}u^{-\frac{((n+1)p-2n)\alpha}{(n+1)p-n}-2}\left\lvert\nabla u\right\rvert^{2p}\eta^{\gamma}\right)^{\frac{p-1}{2p}}$
×$\times\left(\int_{M}u^{-\frac{((n+1)p-2n)\alpha}{(n+1)p-n}+2(p-1)}\left\lvert\nabla\eta\right\rvert^{2p}\eta^{\gamma-2p}\right)^{\frac{1}{2p}},$
@@
$\rho\coloneqq\dfrac{((n+1)-2n)\alpha}{((n+1)p-n)(p-1)\chi_{n,p}}-\dfrac{2}{\chi_{n,p}}>-1,$
@@
Applying Theorem 5.5, we can verify that for each $20<q<$ 2 there holds
@@
[…]
@@
+($+\left(\int_{\left\{u\geq u_{0}\right\}}u^{2(p_{0}-1)-t}\eta^{\gamma}\right)^{\frac{2(p-1)-t}{2(p_{0}-1)-t}}\left(\int_{\left\{u\geq u_{0}\right\}}\left\lvert\nabla\eta\right\rvert^{\frac{(2(p_{0}-1)-t)p}{p_{0}-p}}\eta^{\gamma-\frac{(2(p_{0}-1)-t)p}{p_{0}-p}}\right)^{\frac{2(p_{0}-p)}{2(p_{0}-1)-t}}$
@@
$\int_{\left\{u<u_{0}\right\}}u^{2(p-1)-t}\left\lvert\nabla\eta\right\rvert^{2p}\eta^{\gamma-2p}\leq$ ($\left(\int_{\left\{u<u_{0}\right\}}u^{2\alpha-t}\eta^{\gamma}\right)^{\frac{2(p-1)-t}{2\alpha-t}}\left(\int_{\left\{u<u_{0}\right\}}\left\lvert\nabla\eta\right\rvert^{\frac{(2\alpha-t)p}{\alpha+1-p}}\eta^{\gamma-\frac{(2\alpha-t)p}{\alpha+1-p}}\right)^{\frac{2(\alpha+1-p)}{2\alpha-t}}$

### v3

@@ 4 Gradient estimates
Let $M^{n}$ be a complete and $n$-dimensional Riemannian manifold with the Ricci curvature bounded from below by a nonpositive constant $(n-1)\kappa$. Assume for some constant $\delta_{0}<1$,
@@
Let $M^{n}$ be a complete and $n$-dimensional Riemannian manifold with the Ricci curvature bounded from below by a nonpositive constant $(n-1)\kappa$. Assume $p>1$ and the condition (2.3) holds. If $u$ is a positive and weak solution to (2.1), then the following pointwise inequality holds for every $\lambda\geq\lambda_{0}\coloneqq\frac{p}{1-\delta_{0}^{+}}$
@@
It is well known that $u$ is $C^{\infty}$ in $\left\{\nabla u\neq 0\right\}$ (see [18, 57]). We consider new functions
@@
we have
$\left\langle A_{w}\left(\nabla F\right),\nabla F\right\rangle=p^{2}(p-1)w_{1}^{2p-2}\left((p-1)w_{11}^{2}+\sum_{j=2}^{n}w_{1j}^{2}\right),$ (4.4)
@@
$\left\langle A_{w}\left(\nabla F\right),\mathbf{W}\right\rangle=p(p-1)w_{1}^{2p-2}w_{11}.$ (4.5)
It follows
@@
$-\dfrac{(n-3)p+2}{(n-1)p}\left\langle A_{w}\left(\nabla F\right),\mathbf{W}\right\rangle+\left(\dfrac{(n+1)(p-1)}{n-1}u^{1-p}f(u)-u^{2-p}f^{\prime}(u)\right)F-(n-1)\kappa F^{2-2/p}.$
By (4.4) and (4.5), we have
$|\langle A_{w}(\nabla F),\mathbf{W}\rangle|\leq\sqrt{p-1}F^{1-1/p}\sqrt{\langle A_{w}(\nabla F),\nabla F\rangle}.$
Hence, it follows from the above two inequalities that
$\begin{split}\dfrac{1}{p}\mathcal{L}_{p,w}F\geq&\dfrac{(p-1)^{2}F^{2}}{n-1}+\dfrac{u^{2-2p}f(u)^{2}}{n-1}-\dfrac{2\sqrt{p-1}}{(n-1)p}u^{1-p}\left\lvert f(u)\right\rvert F^{-1/p}\sqrt{\left\langle A_{w}\left(\nabla F\right),\nabla F\right\rangle}\\ &+\dfrac{n(p-1)}{(n-1)p^{2}}F^{-2/p}\left\langle A_{w}\left(\nabla F\right),\nabla F\right\rangle-\dfrac{\left\lvert(n-3)p+2\right\rvert\sqrt{p-1}}{(n-1)p}F^{1-1/p}\sqrt{\left\langle A_{w}\left(\nabla F\right),\nabla F\right\rangle}\\ &+\left(\dfrac{(n+1)(p-1)}{n-1}u^{1-p}f(u)-u^{2-p}f^{\prime}(u)\right)F-(n-1)\kappa F^{2-2/p}.\end{split}$ (4.6)
We note that for any $\delta\in[0,1)$ there holds
$\delta\dfrac{(p-1)^{2}F^{2}}{n-1}+\delta\dfrac{u^{2-2p}f(u)^{2}}{n-1}-\dfrac{2\delta(p-1)}{n-1}u^{1-p}\left\lvert f(u)\right\rvert F\geq 0.$
Now, inserting the above inequality into (4.6) we get
$\dfrac{1}{p}\mathcal{L}_{p,w}F\geq$ $\left(1-\delta\right)\dfrac{(p-1)^{2}F^{2}}{n-1}+\left(1-\delta\right)\dfrac{u^{2-2p}f(u)^{2}}{n-1}+\dfrac{2\delta(p-1)}{n-1}u^{1-p}\left\lvert f(u)\right\rvert F$
@@
for every $\delta\in[0,1)$. Hence, noting that there hold for $\delta=\delta_{0}^{+}=\max\left\{0,\delta_{0}\right\}\in[0,1)$
$\dfrac{(n(1-\delta_{0}^{+})-1)(p-1)}{(1-\delta_{0}^{+})(n-1)p^{2}}=\dfrac{((n-1)(1-\delta_{0}^{+})-\delta_{0}^{+})(p-1)}{(1-\delta_{0}^{+})(n-1)p^{2}}\geq-\dfrac{\delta_{0}^{+}}{(1-\delta_{0}^{+})p},$
and
$\dfrac{\left\lvert(n-3)p+2\right\rvert}{(n-1)p}\leq 1,\quad\mbox{for}\,\,p>1,$
we conclude
@@
For every real number $\lambda>0$, from (4.7) we have
@@
Let $M^{n}$ be a complete and $n$-dimensional Riemannian manifold with the Ricci curvature bounded from below by a nonpositive constant $(n-1)\kappa$. Assume for some constant $\delta_{0}<1$,
@@
Here we have used the following well known fact about the regularity of $u$ (see [57] or [1]):
$u\in C_{loc}^{\infty}\left(M\setminus Z\right)\cap W^{2,2}_{loc}\left(M\setminus Z\right)\cap C^{1,\beta}_{loc}\left(M\right),$
where $Z=\left\{x\in M:\nabla u=0\right\}$ is the critical point set of $u$ on $M$, and for each compact domain $\Omega\Subset M$
$\left\lvert\nabla u\right\rvert^{p-2}\nabla^{2}u\in L^{2}\left(\Omega\setminus\Omega_{cr}\right)$
where $\Omega_{cr}=\left\{x\in\Omega:\nabla u=0\right\}$ is the set of critical points of $u$ on $\Omega$.
@@
On the other hand,
@@
$-\dfrac{4}{p+\lambda_{0}+t-2}\int_{M}\left\langle A_{w}\left(H^{\frac{p+\lambda_{0}+t-2}{2}}\right),\nabla\eta\right\rangle H^{\frac{p+\lambda_{0}+t-2}{2}}\eta.$
Integration by parts, we get
$\dfrac{1}{\lambda_{0}}\int_{M}\eta^{2}H^{t}\mathcal{L}_{p,w}H^{\lambda_{0}}=$ $-\dfrac{4t}{\left(p+\lambda_{0}+t-2\right)^{2}}\int_{M}\left\langle A_{w}\left(\nabla H^{\frac{p+\lambda_{0}+t-2}{2}}\right),\nabla H^{\frac{p+\lambda_{0}+t-2}{2}}\right\rangle\eta^{2}$
@@
$\left(\dfrac{t+1}{t},\,t+1\right),$
@@
$\leq$ $\dfrac{2p(t+1)^{2}}{t}\left(\int_{B_{R}}H^{2t+2}\eta^{2t+2}\right)^{\frac{t}{t+1}}\left(\int_{B_{R}}\left\lvert\nabla\eta\right\rvert^{2t+2}\right)^{\frac{1}{t+1}}+2(n-1)\kappa\left(\int_{B_{R}}H^{2t+2}\eta^{2t+2}\right)^{\frac{t}{t+1}}\left(\int_{B_{R}}\eta^{2t+2}\right)^{\frac{1}{t+1}},$
@@
holds for any cutoff function $\eta$ supported in $B_{R}$. Fixed $q_{0}$ as in 4.2 and set
@@
[…]
@@
Since the Ricci curvature is bounded from below by $(n-1)\kappa$, applying the Bishop-Gromov volume comparison theorem, we get
@@ 5 Liouville theorems
We begin with the following simple case, i.e., $M^{n}$ is a compact manifold without boundary.
@@
Let $u$ be a positive and weak solution to (2.1). By (3.9) we know that the following inequality holds true a.e. on $M^{n}$
$\mathrm{div}\left(u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\nabla_{\mathbf{X}}\mathbf{X}-\dfrac{\mathrm{div}\mathbf{X}}{n}\mathbf{X}\right)\right)\geq 0.$
On the other hand, by Stokes theorem we have
$\int_{M^{n}}\mathrm{div}\left(u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\nabla_{\mathbf{X}}\mathbf{X}-\dfrac{\mathrm{div}\mathbf{X}}{n}\mathbf{X}\right)\right)=0,$
since $M^{n}$ is a compact manifold without boundary. This implies
$\mathrm{div}\left(u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\nabla_{\mathbf{X}}\mathbf{X}-\dfrac{\mathrm{div}\mathbf{X}}{n}\mathbf{X}\right)\right)=0,\quad\mbox{a.e. on}\,\,M^{n}.$
It is easy to see from (3.9) that the above equality holds only if
$\alpha\left(1-\dfrac{(n-p)\alpha}{(n+1)p-n}\right)=0,\quad f(u)=Au^{\alpha},\quad Ric\left(\nabla u,\nabla u\right)=0,\quad\mathrm{tr}\left(\mathscr{X}^{2}\right)=0,$
whenever $\nabla u\neq 0$ and $A$ is a constant. This implies $\nabla u=0$ almost everywhere on $M^{n}$. Thus, $u$ must be a constant function. ∎
Another application of 3.6 is that we provide a new and simple proof of Bidaut-Veron and Veron’s Liouville result [6, Theorem 6.1] on the following equation
$\Delta u-\lambda u+u^{q}=0$
defined on a compact manifold without boundary.
Theorem 5.2 ([6, Theorem 6.1]).
Let $(M,g)$ be a closed manifold with dimension $n\geq 2$ and $u$ be a positive and weak solution to
@@
$0=$ $\int_{M}u^{a}\mathrm{tr}\left(\mathscr{X}^{2}\right)+u^{a+2b}\mathrm{Ric}\left(\nabla u,\nabla u\right)$
@@
$0=$ $\int_{M}u^{\frac{2(n-1)}{n+2}q}\mathrm{tr}\left(\mathscr{X}^{2}\right)+\int_{M}u^{\frac{2(n-3)}{n+2}q}\left(\mathrm{Ric}\left(\nabla u,\nabla u\right)-\dfrac{n-1}{n}\left(q-1\right)\lambda\left\lvert\nabla u\right\rvert^{2}\right)$
$+\int_{M}\frac{n-1}{n+2}q\left(1-\frac{n-2}{n+2}q\right)u^{\frac{2(n-3)}{n+2}q-2}\left\lvert\nabla u\right\rvert^{4}.$
@@ 1. Case 1)
    If $q<\frac{n+2}{n-2}$, then it is easy to see that 0≥∫_M n-1 n+2q(1-n-2n+2q) u^2(n-3)n+2q-2|∇u|^4 implies $\nabla u=0$ and $u$ is a constant.
@@ 2. Case 2)
    If $q=\frac{n+2}{n-2}$, $(M,g)=(\mathbb{S}^{n},g_{0})$ and (5.1) is a strict inequality, then 0≥∫_Mu^2(n-3)n-2 (Ric(∇u,∇u) -n-1n(q-1)λ|∇u|^2) implies $\nabla u=0$ and hence $u$ is a constant.
@@ 3. Case 3)
    If $q=\frac{n+2}{n-2}$, $(M,g)\neq(\mathbb{S}^{n},g_{0})$ and (5.1) is not strict, then we have ∫_Mu^2(n-1)n-2 tr(X^2)=0. Hence, the fact $u>0$ forces $\mathrm{tr}\left(\mathscr{X}^{2}\right)=0$, i.e., tr(X-trXn)^2=0. Noting X=∇(u^b∇u) = ∇(u^-2n-2∇u) = {n-2n-4∇∇(un-4n-2) , if n≠4;∇∇lnu , if n=4, hence, we have
    $\nabla\nabla\left(u^{\frac{n-4}{n-2}}\right)=\frac{1}{n}\Delta\left(u^{\frac{n-4}{n-2}}\right)g,\quad\text{ if }n\neq 4;$
@@
    According to the rigidity theorem [68, Theorem E] (see [35] for more detailed discussions), the equation ∇∇f=1nΔf g admits a non-trivial solution if and only if $(M,g)=(\mathbb{S}^{n},g_{0})$. Thus $u$ must be a constant.
@@
It follows from (3.9) in 3.6, the following inequality holds
$\begin{split}\operatorname{div}\left(u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\mathscr{X}(\mathbf{X})\right)\geq\,&u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\mathrm{tr}\left(\mathscr{X}^{2}\right)\right)+A_{n,p,\alpha}u^{-\frac{((n+1)p-2n)\alpha}{(n+1)p-n}-2}\left\lvert\nabla u\right\rvert^{2p}\\ &+\dfrac{n-1}{n}\left(\alpha f(u)-uf^{\prime}(u)\right)u^{-\frac{((n+1)p-2n)\alpha}{(n+1)p-n}-1}\left\lvert\nabla u\right\rvert^{p}\end{split}$ (5.4)
@@
Direct computation shows
$\int_{M}\operatorname{div}\left(u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\mathscr{X}(\mathbf{X})\right)\eta^{\gamma}=$ $-\gamma\int_{M}u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left\langle\mathscr{X}(\mathbf{X}),\nabla\eta\right\rangle\eta^{\gamma-1}\leq\gamma\int_{M}u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left\lvert\mathscr{X}(\mathbf{X})\right\rvert\left\lvert\nabla\eta\right\rvert\eta^{\gamma-1}$
@@
It follows
$\int_{M}\operatorname{div}\left(u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\mathscr{X}(\mathbf{X})\right)\eta^{\gamma}\leq$ $\sqrt{\frac{1+(p-1)^{2}}{2(p-1)}}\gamma\int_{M}u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\mathrm{tr}\left(\mathscr{X}^{2}\right)\right)^{1/2}u^{-\frac{n(p-1)\alpha}{(n+1)p-n}}\left\lvert\nabla u\right\rvert^{p-1}\left\lvert\nabla\eta\right\rvert\eta^{\gamma-1}$
@@
$\int_{M}\operatorname{div}\left(u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\mathscr{X}(\mathbf{X})\right)\eta^{\gamma}\leq$ $\sqrt{\frac{1+(p-1)^{2}}{2(p-1)}}\gamma\left(\int_{M}u^{\frac{(n-1)p\alpha}{(n+1)p-n}}\left(\mathrm{tr}\left(\mathscr{X}^{2}\right)\right)\eta^{\gamma}\right)^{1/2}\left(\int_{M}u^{-\frac{((n+1)p-2n)\alpha}{(n+1)p-n}-2}\left\lvert\nabla u\right\rvert^{2p}\eta^{\gamma}\right)^{\frac{p-1}{2p}}$
$\times\left(\int_{M}u^{-\frac{((n+1)p-2n)\alpha}{(n+1)p-n}+2(p-1)}\left\lvert\nabla\eta\right\rvert^{2p}\eta^{\gamma-2p}\right)^{\frac{1}{2p}},$
@@
$\rho\coloneqq\dfrac{((n+1)p-2n)\alpha}{((n+1)p-n)(p-1)\chi_{n,p}}-\dfrac{2}{\chi_{n,p}}>-1,$
@@
Applying Theorem 5.5, we can verify that for each $0<q<2$ there holds
@@
Hence, it follows $2p>n$ and we can take $q\in\left(\frac{n}{p},\,2\right)$ such that
@@
$+\left(\int_{\left\{u\geq u_{0}\right\}}u^{2(p_{0}-1)-t}\eta^{\gamma}\right)^{\frac{2(p-1)-t}{2(p_{0}-1)-t}}\left(\int_{\left\{u\geq u_{0}\right\}}\left\lvert\nabla\eta\right\rvert^{\frac{(2(p_{0}-1)-t)p}{p_{0}-p}}\eta^{\gamma-\frac{(2(p_{0}-1)-t)p}{p_{0}-p}}\right)^{\frac{2(p_{0}-p)}{2(p_{0}-1)-t}}$
@@
$\int_{\left\{u<u_{0}\right\}}u^{2(p-1)-t}\left\lvert\nabla\eta\right\rvert^{2p}\eta^{\gamma-2p}\leq$ $\left(\int_{\left\{u<u_{0}\right\}}u^{2\alpha-t}\eta^{\gamma}\right)^{\frac{2(p-1)-t}{2\alpha-t}}\left(\int_{\left\{u<u_{0}\right\}}\left\lvert\nabla\eta\right\rvert^{\frac{(2\alpha-t)p}{\alpha+1-p}}\eta^{\gamma-\frac{(2\alpha-t)p}{\alpha+1-p}}\right)^{\frac{2(\alpha+1-p)}{2\alpha-t}}$